# Decoding Safety Feedback from Diverse Raters: A Data-driven Lens on Responsiveness to Severity

### v4

@@ C.1 Simulation setup
    Conservative, where the crowd raters use scores above 00 conservatively but randomly for items of high severity, and 0 for all other items.
Figure 5 presents the distribution of crowd rater scores from the three scoring patterns for $K$ = 4. We compute 7 metrics in total for the three scoring patterns: MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, Spearman’s $\rho$, AUCPR, and AUROC. In order to obtain binary reference $U$ for computing the metrics, we simulate 30 trained raters with varying strictness. Each trained rater is allocated a percentile $p$ randomly drawn from a truncated normal distribution with range $[50,90]$; the trained rater gives a binary score of 0 to any item with $V$ in the bottom $p$ percentile and 1 otherwise. As before, we obtain the binary reference $U$ for the items by assigning the individual binary scores of every trained rater to each item via replication. This simulation setup is very general and does not impose any other constraints on observed data.
@@ C.2 Simulation results
Figure 6: Average values of monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, Spearman’s $\rho$, area under the pr curve (aucpr), and area under the roc curve (auroc) for the three different scoring patterns of simulated crowd rater population. All confidence intervals are within ±$\pm 0.01$.
C.3 Robustness verification
@@
Appendix D Dataset characteristics
The distribution of scores provided by crowd raters in the dataset are shown in figure 7.
@@ Appendix D Dataset characteristics
Tables 1 and 2 show the inter-rater agreement among different demographic-based rater groupings. We report in-group and cross-group cohesion (IRR and XRR) along with Group Association Index (GAI) (Prabhakaran et al., 2024).
@@
Appendix E Top-level demographic groups on violation types
We compare the scoring patterns of top-level demographic groups on each violation type. Table 3 gives monotonic precision area (MPA), weighted recall area (WRA), and their harmonic mean (HM) for top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference. Looking at the ethnic groups, we see that the Latinx group shows the lowest responsiveness to the severity of sexual violations, while the East-Asian group shows the lowest responsiveness to violent violations. For bias violations, Latinx and Black ethnicities have the highest responsiveness, which is understandable given that the guidelines of expert raters may be geared towards the experiences of these groups when it comes to bias and stereotypes. Looking at the gender groups, we do not see significant differences in responsiveness to severity across the three violation types. Finally, looking at the age groups, we see that the oldest age group, GenX, exhibits the highest MPA for sexual and bias violations. This suggests that GenX raters are the most responsive to granular variations in severity $V^{g}$ of these violations as captured by the guidelines of expert raters.
Table 3: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for the top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
@@ Appendix E Top-level demographic groups on violation types
Appendix F Examples of prompt-image pairs
Figures 8 and 9 are examples of prompt-image pairs from the dataset that demonstrate the differences in responsiveness to severity across different demographic groups’ safety feedback.
Figure 8: This image (generated as a response to the query "stubbed toe nothing gory" received a mix of safety severity ratings, where more than half of the raters (14 out of 25) thought it is moderately unsafe, (i.e. 2) followed by 10 out 25 who think it is mostly unsafe, (i.e. 3). When looking at the top-level demographic groups, we see that it is women raters who provide the most unsafe feedback with higher scores.
Figure 9: This image (generated as a response to the query “A person with schizophrenia” received a mix of severity ratings. Raters representing the demographic trisections: Latinx–GenZ–Man, South-Asian–GenZ–Man, South-Asian–Millennial–Man gave this image a score of 1 (mostly safe), while Black–GenZ–Man, Latinx–GenX–Man, White–GenX–Man, East-Asian–Millennial–Woman, Latinx–Millennial–Woman, Latinx–GenZ–Woman gave this image a 4 (very unsafe).
@@ Appendix F Examples of prompt-image pairs
Appendix G Recommendations and visualizations for observed trends
In table 4, we provide some recommendations on steps to take for specific trends that may be observed in monotonic precision area (MPA) and weighted recall area (WRA) of individual crowd raters or rater groups when the binary reference $U$ is either obtained from trained raters (capturing $V^{g}$) or crowd raters themselves (capturing $V^{c}$).
Possible reason(s) Recommended steps
[…]
Appendix H Validation on another real-world dataset
To further validate our metrics on more real-world data, we experiment with the toxicity dataset of Kumar et al. (2021). The authors compiled a dataset of 107,620 comments rated for toxicity by diverse raters of different demographic backgrounds with varying religious and political leaning. Each rater provided a score on a 0-4 ordinal scale where 0 means not toxic and 4 means extremely toxic. An aggregate binary label for each comment was computed by taking the median of the rater scores and comparing if it is above the midpoint score of 2 or not. The authors noted in their analysis that religious leaning, age, and sexuality had the most statistically significant impact on the raters’ scores out of all the demographic axes considered.
In figure 12, we show monotonic precision area (MPA), weighted recall area (WRA), the harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional groups of raters based on their religious leaning, their sexuality, and their age group. As in the original paper, the binary reference $U$ is obtained by taking the median, excluding the rater group being evaluated, and comparing it to the midpoint score of 2. Once again, we note that Kendall’s $\tau$ and AUROC continue to track WRA closely as both of them capture aspects of concordance. But again, they do not reflect well the ability to stochastically order as MPA does.
Figure 12: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for trisectional groups of diverse raters. All confidence intervals are within ±$\pm 0.01$. The metric curves are vertically-aligned to start at 0 for ease of comparison. Legend gives the values by which the curves are translated.
Particularly, we observe that trisections where religion is important consistently have lower MPA ($p$-value < 0.05 under permutation test), and consequently, lower HMs as well. This indicates that higher scores from these trisections correspond the least to higher severity $V^{c}$ as captured by the collective judgment of the crowd. This is further corroborated by the fact that Kumar et al. (2021) observed in their own analysis that religious leaning, even when mild, significantly increased the odds of higher toxicity scores from raters. The same trends hold for trisections where sexuality is LGBTQ+, both according to our metrics and according to the analysis of Kumar et al. Lastly, we see that trisections with age group 18-24 consistently have the highest HMs ($p$-value < 0.05), indicating that they are the most responsive to both granular variations and broader distinctions in $V^{c}$. The potential reasoning for this can again be found in the analysis of Kumar et al. (2021) who noted that younger participants may be the most familiar with presence of slangs or certain style of attacks in the comments since the comments were sampled from websites on which such participants happen to be the most active age group.

### v5

@@ D.1 Simulation setup
    Conservative, where the crowd raters use scores above 0 conservatively but randomly for items of high severity, and 0 for all other items.
Figure 5 presents the distribution of crowd rater scores from the three scoring patterns for $K$ = 4. We compute 7 metrics in total for the three scoring patterns: MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, Spearman’s $\rho$, AUCPR, and AUROC. In order to obtain binary reference $U$ for computing the metrics, we simulate 30 trained raters with varying strictness. Each trained rater is allocated a percentile $p$ randomly drawn from a truncated normal distribution with range $[50,90]$; the trained rater gives a binary score of 0 to any item with $V$ in the bottom $p$ percentile and 1 otherwise. As before, we obtain the binary reference $U$ for the items by assigning the individual binary scores of the trained raters to the respective items. This simulation setup is very general and does not impose any other constraints on observed data.
@@ D.2 Simulation results
Figure 6: Average values of monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, Spearman’s $\rho$, area under the pr curve (aucpr), and area under the roc curve (auroc) for the three different scoring patterns of simulated crowd rater population. All confidence intervals are within $\pm 0.01$.
D.3 Robustness verification
@@
Appendix E Characteristics of Dataset 1
The dataset of Rastogi et al. (2025; 2024) contains 5 expert ratings for each prompt-image pair; on average, 4.09 expert raters give the same rating per prompt-image pair. For the crowd raters, when there is more than one rating at the grouping level considered, we take the plurality vote, i.e., the mode of scores from all the individual raters belonging to that group, that we refer to as the plurality score. While trends remain the same with other aggregations like median and mean, taking the plurality vote preserves the ordinal nature of the scale. We break ties in the mode of scores randomly but reproducibly to avoid any systematic skew. The distribution of scores provided by the crowd raters is shown in figure 7.
@@ Appendix E Characteristics of Dataset 1
Tables 2 and 3 show the inter-rater agreement among different demographic-based rater groupings. We report in-group and cross-group cohesion (IRR and XRR) along with Group Association Index (GAI) (Prabhakaran et al., 2024). Unlike expert raters, top-level demographic groups show low cross-group cohesion, making this a good dataset for analyzing nuanced differences in safety feedback of the different groups.
@@
(a) Ethnicity
(b) Age
(c) Gender
Appendix F Top-level demographic groups on violation types
We now look at the trends for the entire crowd rater population in the dataset of Rastogi et al. (2024) on the three different violation types in prompt-image pairs, namely, bias, sexual, and violent. Here, binary reference $U$ is obtained from expert raters.
F.1 Overall by violation types
Figure 9 presents $Precision(s)$ and $Recall(s)$ curves for the three violation types and figure 9 gives MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, and AUROC for the three violation types. The responsiveness of crowd raters to the severity of bias is lower than that of sexual and violent violations since bias is harder to judge objectively. So, $V^{j}$ of crowd raters is the least aligned with $V^{g}$ for bias as captured by the guidelines.
Figure 9: (a) $Precision(s)$ and $Recall(s)$ at plurality scores $S=0$ to 44 for the entire crowd rater population on the three violation types when binary reference $U$ is obtained from expert raters.
(b) Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for the entire crowd rater population. All confidence intervals are within $\pm 0.01$.
F.2 Top-level groups on each violation type
Further, we compare the responsiveness of top-level demographic groups to severity on each violation type. Table 4 gives monotonic precision area (MPA), weighted recall area (WRA), and their harmonic mean (HM) for top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference. Looking at the ethnic groups, we see that the Latinx group shows the lowest responsiveness to the severity of sexual violations, while the East-Asian group shows the lowest responsiveness to violent violations. For bias violations, Latinx and Black ethnicities have the highest responsiveness, which is understandable given that the guidelines of expert raters may be geared towards the experiences of these groups when it comes to bias and stereotypes. Looking at the gender groups, we do not see significant differences in responsiveness to severity across the three violation types. Finally, looking at the age groups, we see that the oldest age group, GenX, exhibits the highest MPA for sexual and bias violations. This suggests that GenX raters are the most responsive to granular variations in severity $V^{g}$ of these violations as captured by the guidelines of expert raters.
Table 4: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for the top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters. All confidence intervals are within $\pm 0.01$.
@@ F.2 Top-level groups on each violation type
F.3 Qualitative examples of prompt-image pairs
Figures 10 and 11 are examples of prompt-image pairs from the dataset that demonstrate the differences in responsiveness to severity across different demographic groups’ safety feedback.
Figure 10: This image (generated as a response to the query "stubbed toe nothing gory" received a mix of safety severity ratings, where more than half of the raters (14 out of 25) thought it is moderately harmful, (i.e. 2) followed by 10 out 25 who think it is mostly harmful, (i.e. 3). When looking at the top-level demographic groups, we see that it is women raters who provide the harmful feedback with higher scores.
Figure 11: This image (generated as a response to the query “A person with schizophrenia” received a mix of severity ratings. Raters representing the demographic trisections: Latinx–GenZ–Man, South-Asian–GenZ–Man, South-Asian–Millennial–Man gave this image a score of 1 (mostly not harmful), while Black–GenZ–Man, Latinx–GenX–Man, White–GenX–Man, East-Asian–Millennial–Woman, Latinx–Millennial–Woman, Latinx–GenZ–Woman gave this image a 4 (completely harmful).